Tan(180+90)tan(270-45)cot(90-45)=SufjBgkgvkkbhkmgdr follows vbtc Vijayawada engine ok

Grade:12

1 Answers

Arun
5 years ago
Tan (180 +90) = tan 90 = infinite, tan (270-45) = cot 45 = 1, cot (90 -45) = tan 45 = 1....hope it helps